The answer to each of the questions is a single digit integer ranging from 0 to 9.


 






The velocity of water waves (V) may depend on their wavelength l, the density of water r and the acceleration due to gravity, g.



 

Grade:10

1 Answers

Navjyot Kalra
10 years ago

5

 

        V = klapbgc

 

        [M0LT–1] = La (ML–3)b (LT–2)c = Mb La–3b+c T–2c

 

        \ b = 0; a – 3b + c = 1

 

        –2c = –1 ® c = ½                       \ a =

 

        V µ l1/2 p0 g1/2             or     V2 µ l g

 

        \ 2x + 3y = 5